OpenStudy (anonymous):

if every side of a cube is increased 4 times, its volume will increased how many times?

OpenStudy (nubeer):

64 times

OpenStudy (unklerhaukus):

\[V_{\text{cube}}=s^3\] \[V_{\text{big cube}}=(4s)^3=4^3\times s^3\]

OpenStudy (anonymous):

Area increases with the square and volume by the cube

OpenStudy (anonymous):

let the side be 'x' hence the volume will be x^3 if it is 4 times increased, the side will be 4x and the volume will be (4x)^3= 64x^3 if compared, its 64 times increased ;)

OpenStudy (anonymous):

@meme_90 what is the answer?